Calculate the concentration of HNO3 in a solution at 25

°

C that has a pH of 4.07.
Answer #1

HNO3 is a HNO → strong acid Hilay) + NO5 (ay) pra 4.07 10-PH [H+1: = 10-4.07 - 8.51x105m [HINO3)= [ht] 8. 51 x los